can you pls sent me a complete table of formulas in finding the equation of the line?
i mean, given two points, slope, etc..
pls help....

Answer by venugopalramana(3286)

1..SLOPE=M....INTERCEPT=C..........EQN.IS ....Y=MX+C
2..POINT =(X1,Y1)....AND.....SLOPE=M..EQN.IS....Y-Y1=M(X-X1)
3..2 POINTS...=(X1,Y1) AND (X2,Y2)..EQN.IS...(Y-Y1)/(Y2-Y1)=(X-X1)/(X2-X1)
4..2 INTERCEPTS..X INTERCEPT=A...Y INTERCEPT=B.....EQN.IS....(X/A)+(Y/B)=1